OpenStudy (foxythepirate0984):

2/3 cups are made to make 6 pancakes, how much cups is it to make 1 pancake

OpenStudy (foxythepirate0984):

@mehek Can you please help?

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} cups&pancakes \\\hline\\ \frac{2}{3}&6\\ x&1 \end{array}\implies \cfrac{\frac{2}{3}}{x}=\cfrac{6}{1}\)

OpenStudy (youngstudier):

Are there answer choices, @Foxythepirate0984 ?

OpenStudy (youngstudier):

@Foxythepirate0984 ? Are there answer choices, Foxy?

OpenStudy (foxythepirate0984):

sorry i was doing another question till someone helps.

OpenStudy (youngstudier):

Foxy? Ish dere answer choices? \(^_^)/

OpenStudy (foxythepirate0984):

no.

OpenStudy (youngstudier):

Drat....... One second then.....

OpenStudy (foxythepirate0984):

second what?

OpenStudy (youngstudier):

Imma google it. :P

OpenStudy (foxythepirate0984):

ok

OpenStudy (youngstudier):

brb

OpenStudy (youngstudier):

Drat, google doth not have it.... Is this a test question or a workbook question @Foxythepirate0984 ?

OpenStudy (foxythepirate0984):

test question

OpenStudy (youngstudier):

I was hoping it was a workbook question, cuz then you could skip it. :P

OpenStudy (foxythepirate0984):

ikr. in still confused what to do or how to solve it..

OpenStudy (youngstudier):

@jdoe0001

OpenStudy (youngstudier):

Wait a second, you've asked this question before.... 6=>2/3 1=>x x/1=(2/3)/6 x=2/3*1/6=2/18=simplify make a ratio ....Let's see if we can figure this out. ^u^

OpenStudy (foxythepirate0984):

i asked it before?

OpenStudy (youngstudier):

yep, the thing I just posted is the closest thing that can help us figure this out.

OpenStudy (youngstudier):

Let's see, 2 over 3 divided by 6 so: 2/3 * 1/6 What does that equal?

OpenStudy (foxythepirate0984):

2/18?

OpenStudy (youngstudier):

Yep.

OpenStudy (youngstudier):

Simplified 2/18 = ?

OpenStudy (youngstudier):

@Foxythepirate0984 ?

OpenStudy (foxythepirate0984):

9

OpenStudy (foxythepirate0984):

@YoungStudier

OpenStudy (foxythepirate0984):

@YoungStudier hello?

OpenStudy (youngstudier):

Sorry, I'm here.

OpenStudy (youngstudier):

Um, no actually. 2/18 2 goes into 2 once. 2 goes into 18 nine times So it's 1 over 9 or 1/9.

OpenStudy (youngstudier):

@Foxythepirate0984 ?

OpenStudy (foxythepirate0984):

thanks ^u^

OpenStudy (youngstudier):

No problem :)

OpenStudy (youngstudier):

Tell me if you got it right. :D

OpenStudy (youngstudier):

@Foxythepirate0984 , how'd you do?

OpenStudy (youngstudier):

Well, I hope you mastered it. :) Talk to you later, then. ^u^